solve by elimination method
0.3x-0.2y=4
1.04x+0.5y=1

Answer by mananth(16946)

multiply first equation by 10
second equation by 100
3 x + 2 y = 40 .............1
104 x + 50 y = 100 .............2
Eliminate y
multiply (1)by -25
Multiply (2) by 1
-75 x -50 y = -1000
104 x + 50 y = 100
Add the two equations
29 x = -900
/ 29
x = -31.03
plug value of x in (1)
3 x + 2 y = 40
-93.1 + 2 y = 40
2 y = 40 + 93.1
2 y = 133.1
y = 66.55
m.ananth@hotmail.ca